OpenStudy (anonymous):

Use implicit differentiation to find the following: 6x^2+2y^2=42 (a) The slope of the tangent line at the indicated point on the graph m=? (b) The equation of the tangent line at the indicated point on the graph. y=?

OpenStudy (anonymous):

slope at m = 12x/4y

OpenStudy (helder_edwin):

\[ \LARGE 12x+4yy'=0\quad\Rightarrow\quad y'=? \]

OpenStudy (anonymous):

forgot - sign

OpenStudy (anonymous):

shoot i forgot to add at point (2,-3)!!!!!

OpenStudy (anonymous):

12*3/3*4=3

OpenStudy (anonymous):

how about part b?

OpenStudy (anonymous):

y+3=3(x-2)

OpenStudy (anonymous):

so the answer is?..

OpenStudy (anonymous):

y=3x-3

OpenStudy (anonymous):

its saying that it is incorrect :(

OpenStudy (anonymous):

ups sry: y=3x-9

OpenStudy (anonymous):

still incorrect :(

OpenStudy (anonymous):

y=2x-7

OpenStudy (anonymous):

thats right thank you!! how about this one, same concept:

OpenStudy (anonymous):

use implict differentiation to find the following: 6x^2-y^2=xy, (-1,3) (a) The slope of the tangent line at the indicated point on the graph. m=? (b) The equation of the tangent line at the indicated point on the graph. y=?

OpenStudy (helder_edwin):

\[ \LARGE y'=-\frac{12x}{4y}=-\frac{3x}{y} \] \[ \LARGE y'|_{(2,-3)}=-\frac{3\cdot 2}{-3}=2 \] the tangent line has equation \[ \LARGE y-y_0=y'(x_0,y_0)\cdot(x-x_0) \] so \[ \LARGE y-(-3)=2(x-2) \] \[ \LARGE y=2x-7 \]

OpenStudy (helder_edwin):

you should be able to do the other problem on you own!

OpenStudy (anonymous):

thank you helder!

OpenStudy (anonymous):

i got that m=-3 but not the y=

OpenStudy (anonymous):

could anyone help me with that part